# Offline changepoint localization using a matrix of conformal p-values

### v4

@@ 6.2 Sentiment change using large language models (LLMs)
Next, we consider a simulation with a sentiment change in a sequence of text samples, showing that our algorithm is practical for localizing changepoints in language data. We consider the Stanford Sentiment Treebank (SST-2) dataset of movie reviews labeled with a binary sentiment, introduced in [24]. Suppose that we would like to localize a change from a generally positive sentiment in movie reviews to a negative sentiment. In practice, this method may be applied to localize a change in customer sentiment toward a product or general approval of a political leader.
@@
As described in Section 5, we can use a pretrained multi-class classifier to estimate the likelihood ratio using the data before and after $t$ (for each $t\in[n-1]$). For this simulation, we use the DistilBERT base model, fine-tuned for sentiment analysis on the uncased SST-2 dataset ([20]).
@@
Figure 5: p-values for SST-2 sentiment change at $\xi=400$ using DistilBERT trained for sentiment analysis. The dashed red line indicates the true changepoint, and the region on the horizontal axis where the p-values lie above the dotted green line ($\alpha=0.05$) corresponds to our 95% confidence set. The point estimators are a) $\hat{\xi}=388$ and b) $\hat{\xi}=373$, both of which are close to the true $\xi=400$.
@@ 6.3 Human activity change: accelerometer data
Here, we include a simulation using data from the Human Activity Recognition (HAR) dataset, introduced in [3]. This dataset contains tri-axial accelerometer data collected from a smartphone accelerometer mounted at the waist of 30 human subjects during various activities. Suppose we observe $n=1000$ samples accelerometer with a changepoint at $\xi=400$; the pre-change class $\mathbb{P}_{0}$ consists of accelerometer samples from the “walking upstairs” activity and the post-change class $\mathbb{P}_{1}$ consists of samples from the “standing” activity. For simplicity, we only use accelerometer data from the first axial direction, so the data we work with is real-valued. We show the dataset in Figure 6.
@@
Then, as described in Section 5, we can use a kernel density estimator (KDE) to estimate the likelihood ratio using the data before and after $t$ (for each $t\in[n-1]$), using a Gaussian kernel. We choose bandwidth $10^{-1}$ when $t\in\{1,n-1\}$ and bandwidth $r^{-1/5}$ otherwise, where $r$ is the number of samples used to learn each KDE; this is known as Scott’s rule (due to [21]). However, due to dependence of the left and right p-values, we use the Bonferroni combining rule, plotting the resulting p-values $(p_{t})_{t=1}^{n-1}$ in Figure 7.
The resulting confidence set is $[376,444]$ and contains the changepoint, as shown in Figure 7. Note that we were able to obtain such a tight confidence interval without any assumptions about the nature of the change. As discussed in Section 2, most prior nonparametric changepoint localization focus on localizing a change in mean. However, for our dataset, the change in mean between the left and right halves is only 0.003-0.003, making it difficult to localize using standard methods.
@@
Finally, consider a simulation of a digit change from the CIFAR-100 image dataset [12]; even when the data is high-dimensional with complex structure, our algorithm can be used to localize the changepoint. This dataset contains $32\times 32\times 3$ RGB images from 100 classes, making the localization problem challenging. Suppose that we observe data of length $n=800$ and the changepoint is $\xi=300$; the pre-change class $\mathbb{P}_{0}$ consists of i.i.d. draws from the set of bear images and the post-change class $\mathbb{P}_{1}$ consists of i.i.d. draws from the set of beaver images (Figure 8). Note that bears and beavers are visually difficult to tell apart, so localization of the changepoint is a difficult task.
@@
We can then use the MCP algorithm and Algorithm 2 to produce a confidence interval. In this case, the left and right p-values are independent, since the left score function $s_{t}^{(0)}$ only depends on data to the left of $t$ and the right score function $s_{t}^{(1)}$ only depends on data to the right of $t$. Therefore, we use the minimum method to combine the p-values, plotting the resulting p-values $(p_{t})_{t=1}^{n-1}$ in Figure 9; the resulting confidence set is $[263,316]$.
@@ Acknowledgments
We thank Carlos Padilla for early discussions, Jing Lei for discussions relating to the conformal Neyman-Pearson lemma, Swapnaneel Bhattacharyya for noticing an error in an early preprint, and Alessandro Rinaldo for several references to prior work in changepoint localization.
Appendix A Gaussian mean change
Even if we didn’t know that the data contained a changepoint, we can pre-test the whole data for exchangeability, as discussed in Appendix D. When we pre-test the data for exchangeability at level $\alpha_{0}=0.01$ using the identity score function and only the forward p-values, we estimate over 1000 simulations (with no changepoint) that the Type I error rate of our pre-testing algorithm is $0.009$, which aligns closely with the theoretical Type I error of $\alpha_{0}=0.01$. Over 1000 simulations (with a changepoint), the pre-test was always able to detect a deviation from exchangeability, and had an empirical power of 1.000.
For all future simulations, assume that we do not use a pre-test for exchangeability. Suppose that we do not have access to the oracle score function. Then, as described in Section 5, we can use a kernel density estimator to estimate the likelihood ratio using the data before and after $t$ (for each $t\in[n-1]$), using a Gaussian kernel. We choose bandwidth $10^{-1}$ when $t\in\{1,n-1\}$ and bandwidth $r^{-1/5}$ otherwise, where $r$ is the number of samples used to learn each KDE; this is known as Scott’s rule (due to [21]). Then, we repeat the process described above to compute the left and right p-values. However, due to dependence of the left and right p-values, we use the Bonferroni combining rule, plotting the resulting p-values $(p_{t})_{t=1}^{n-1}$ in Figure 10b.
(a) oracle score function family
(b) KDE learned score function family
Figure 10: p-values for a Gaussian mean change at $\xi=400$. The dashed red line indicates the true changepoint, and the region on the horizontal axis where the p-values lie above the dotted green line ($\alpha=0.05$) corresponds to our 95% confidence set. The point estimators are a) $\hat{\xi}=409$ and b) any $\hat{\xi}=380$, both of which are close to the true $\xi=400$.
To give more intuition for the MCP algorithm, we can also visualize the matrix of conformal p-values (MCP) as described in Section 3, in Figure 11. Observe that only when $t=\xi=400$ are the p-values in row $t$ truly uniformly distributed and independent in the left and right halves respectively, as they should be under $\mathcal{H}_{0t}$. For instance, if $t<\xi$, then the right p-values will be invalid and the Kolmogorov-Smirnov statistic $W_{t}^{(1)}$ used in the MCP algorithm (Algorithm 1) will detect their deviation from uniformity.
(a) matrix of conformal p-values (MCP)
(b) validity of conformal p-values
Running the same simulations for 1000 trials, we report the empirical average width and coverage of our confidence sets at various levels, as well as the bias and mean absolute deviation of our point estimator $\hat{\xi}=\arg\max_{t\in[n-1]}\;p_{t}$, in Table 1. In Table 1, we also compare the performance of the oracle and (KDE) learned scores against a parametric learned score. The parametric learned score is similar to the KDE learned score, except that we only learn the mean and assume that our data is Gaussian with unit variance when computing the likelihood ratio.
Method Avg. width Coverage Bias Mean absolute deviation
Oracle score (50%) 22.56 $0.50_{\pm 0.03}$ -1.01 13.13
Parametric learned score (50%) 23.74 $0.51_{\pm 0.03}$ 0.20 13.10
KDE learned score (50%) 24.08 $0.52_{\pm 0.03}$ 12.13 35.09
Oracle score (95%) 74.33 $0.94_{\pm 0.02}$ -1.01 13.13
Parametric learned score (95%) 75.34 $0.95_{\pm 0.02}$ 0.19 13.04
KDE learned score (95%) 75.82 $0.97_{\pm 0.02}$ 3.46 15.29
Table 1: Average width and coverage of confidence sets for Gaussian mean change, as well as the bias and mean absolute deviation of the point estimator.
We know from Theorem 3.1 that the true coverage probability is exactly 50% and 95% respectively for a 50% or 95% confidence interval constructed using our method; this is what we observe empirically as well. Importantly, the average width of the confidence sets using the learned score function is very comparable to the average width of the confidence sets from using the oracle score function, showing that our method performs well even when the likelihood of the data is not available. Furthermore, though we do not give theoretical guarantees for our point estimator, we observe that the point estimator is empirically close to the true changepoint.
@@ Appendix B Using certified data to tighten confidence sets
Consider the MNIST digit change experiment of Section 6 but suppose we had additional certified pre-change and post-change samples, which we add to the left and right samples respectively. In this case, we are able to obtain better results.
@@
Figure 12: p-values for an MNIST digit change at $\xi=400$ using 100 certified calibration samples (a) on the left side only and (b) on each side. The dashed red line indicates the true changepoint, and the region on the horizontal axis where the p-values lie above the dotted green line ($\alpha=0.05$) corresponds to our 95% confidence set. The point estimators are a) $\hat{\xi}=393$ and b) $\hat{\xi}=405$, both of which are close to the true $\xi=400$.
@@
Appendix C Extensions of the MCP algorithm
@@ C.1 Alternative tests
We can use the asymptotic distribution of $W_{t}$ to construct confidence sets for the changepoint. This method is valid when $t$ and ntn-t are both large. Recall that $\varphi_{t}=B_{t}-tB_{1}$ (where $(B_{t})_{t\geq 0}$ is a standard Brownian motion on $\mathbb{R}$) is the Brownian bridge on $\mathbb{R}$. Define $F_{\varphi}(z):=\mathbb{P}\left(\sup_{t\in[0,1]}\;\lvert\varphi_{t}\rvert\leq z\right)$, where $(\varphi_{t})_{t\geq 0}$ is the Brownian bridge. Then, by a central limit theorem due to [7], we can compute the asymptotically valid p-values
@@
Note that in general, one can use the approximation $p_{t}^{\mathrm{left}}\approx 2\exp(-2\,(W_{t}^{(0)})^{2})$, which is obtained by truncating the series expansion of the limiting distribution ([11]). In practice, we can use the empirical test to test $\mathcal{H}_{0t}$ when $t$ and ntn-t are small, and use the asymptotic test when $t$ and ntn-t are large. Next, we will discuss how to combine the left and right p-values from Section 4.1 into a single p-value for $\mathcal{H}_{0t}$.
@@
Theorem C.1 (Asymptotic coverage guarantee (asymptotic KS test)).
@@
Input: $(W_{t}^{(0)},W_{t}^{(1)})$ (discrepancy scores), $B\in\mathbb{N}$ (sample size)
@@
1
2for $b\in[B]$ do
3 $(X_{t}^{(b)})_{t=1}^{n}\sim\mathrm{Unif}([0,1]^{n})$
4  Compute $(W_{t}^{(0,b)})_{t=1}^{n-1}$ and $(W_{t}^{(1,b)})_{t=1}^{n-1}$ using Algorithm 1 to Algorithm 1 of Algorithm 1 on permuted data $(X_{t}^{(b)})_{t=1}^{n}$
5 end for
6
7Draw $\theta_{0},\,\theta_{1}\sim\mathrm{Unif}(0,1)$
[…]
@@ 3 $(X_{t}^{(b)})_{t=1}^{n}\sim\mathrm{Unif}([0,1]^{n})$
C.2 Alternatives to the Kolmogorov-Smirnov statistic
@@
Appendix D Pre-testing for exchangeability
Suppose we are unsure whether a changepoint exists in the dataset. In this case, it is natural to first test the entire data for exchangeability and only proceed with the MCP algorithm if we conclude that the data is not exchangeable. Formally, we are testing
$\mathcal{H}_{0}:(X_{1},\dots,X_{n})\text{ are exchangeable}$
against
$\mathcal{H}_{1}:(X_{1},\dots,X_{n})\text{ are not exchangeable}.$
Following [28], using a score function $s$, we compute the sequential (randomized) ranks for $t\in[n]$ as
$\tilde{p}_{t}=\frac{1}{t}\sum_{j=1}^{t}(\mathbf{1}_{s(X_{j})>s(X_{t})}+\theta_{t}\,\mathbf{1}_{s(X_{j})=s(X_{t})}),$
where $\theta_{t}\sim\mathrm{Unif}(0,1)$ are i.i.d. and independent of the data. Under exchangeability, $p_{t}\sim\mathrm{Unif}(0,1)$ are independent, so we can perform a one-sample KS test from uniform using $(\tilde{p}_{1},\dots,\tilde{p}_{n})$. In particular, if $\hat{F}$ is the empirical cdf of $(\tilde{p}_{1},\dots,\tilde{p}_{n})$ and $u$ is the cdf of $\mathrm{Unif}(0,1)$, we can compute the statistic $\sqrt{n}\;\mathrm{KS}(\hat{F},u)$ and reject when this statistic is above a threshold $t_{\alpha_{0}}$, which is chosen so that the Type I error of the test does not exceed $\alpha_{0}\in(0,1)$. Here, the threshold can be chosen by empirically simulating uniform random variables, as in Section 4.1. The p-value associated with this test is termed the “forward p-value”.
We also recommend to compute the sequential ranks in reverse by
and calculate a “backward p-value” use a one-sample KS test from uniform using $(\overline{p}_{1},\dots,\overline{p}_{n})$. Then, the forward and backward p-values obtained from the respective KS tests can be combined using the Bonferroni method (rejecting only when either of the p-values is at least $2\alpha_{0}$).
Hence, it suffices to choose a good score function to test exchangeability, for which several methods are provided in [22]. If we pre-test the data for exchangeability using a test of level $\alpha_{0}\in(0,1)$, then our CI algorithm will lose its theoretical guarantees on the set (of probability at most $\alpha_{0}$) where the pre-test makes a Type I error, but we show in Appendix A that this does not affect the empirical performance of the MCP algorithm. Next, we discuss how the score function can be learned from the data to tighten our confidence sets.
Appendix E Proofs of main results
In this section, we present proofs of our main results. For an alternative proof technique, see [29, 2]. We begin with a useful lemma (for more on this lemma, see [4]).
Lemma E.1 (Randomized probability integral transform).

### v5

@@ 7.2 Sentiment change using large language models (LLMs)
Next, we consider a simulation with a sentiment change in a sequence of text samples, showing that our algorithm is practical for localizing changepoints in language data. We consider the Stanford Sentiment Treebank (SST-2) dataset of movie reviews labeled with a binary sentiment, introduced in [socher2013recursive]. Suppose that we would like to localize a change from a generally positive sentiment in movie reviews to a negative sentiment. In practice, this method may be applied to localize a change in customer sentiment toward a product or general approval of a political leader.
@@
As described in Section 5, we can use a pretrained multi-class classifier to estimate the likelihood ratio using the data before and after $t$ (for each $t\in[n-1]$). For this simulation, we use the DistilBERT base model, fine-tuned for sentiment analysis on the uncased SST-2 dataset ([sanh2019distilbert]).
@@
Figure 9: p-values for SST-2 sentiment change at $\xi=400$ using DistilBERT trained for sentiment analysis. The dashed red line indicates the true changepoint, and the region on the horizontal axis where the p-values lie above the dotted green line ($\alpha=0.05$) corresponds to our 95% confidence set. The point estimators are (a) $\hat{\xi}=388$ and (b) $\hat{\xi}=373$, both of which are close to the true $\xi=400$.
@@ 7.3 Human activity change: accelerometer data
Here, we include a simulation using data from the Human Activity Recognition (HAR) dataset, introduced in [anguita2013public]. This dataset contains tri-axial accelerometer data collected from a smartphone accelerometer mounted at the waist of 30 human subjects during various activities. Suppose we observe $n=1000$ samples accelerometer with a changepoint at $\xi=400$; the pre-change class $\mathbb{P}_{0}$ consists of accelerometer samples from the “walking upstairs” activity and the post-change class $\mathbb{P}_{1}$ consists of samples from the “standing” activity. For simplicity, we only use accelerometer data from the first axial direction, so the data we work with is real-valued. We show the dataset in Figure 10.
@@
Then, as described in Section 5, we can use a kernel density estimator (KDE) to estimate the likelihood ratio using the data before and after $t$ (for each $t\in[n-1]$), using a Gaussian kernel. We choose bandwidth $10^{-1}$ when $t\in\{1,n-1\}$ and bandwidth $r^{-1/5}$ otherwise, where $r$ is the number of samples used to learn each KDE; this is known as Scott’s rule (due to [scott1979optimal]). However, due to dependence of the left and right p-values, we use the Bonferroni combining rule, plotting the resulting p-values $(p_{t})_{t=1}^{n-1}$ in Figure 11.
The resulting confidence set is $[376,444]$ and contains the changepoint, as shown in Figure 11. Note that we were able to obtain such a tight confidence interval without any assumptions about the nature of the change. As discussed in Section 2, a lot of work on prior nonparametric changepoint localization focuses on localizing a change in mean. However, for our dataset, the change in mean between the left and right halves is only 0.003-0.003, making it difficult to localize using this class of methods.
@@
Finally, consider a simulation of a digit change from the CIFAR-100 image dataset [krizhevsky2009learning]; even when the data is high-dimensional with complex structure, our algorithm can be used to localize the changepoint. This dataset contains $32\times 32\times 3$ RGB images from 100 classes, making the localization problem challenging. Suppose that we observe data of length $n=800$ and the changepoint is $\xi=300$; the pre-change class $\mathbb{P}_{0}$ consists of i.i.d. draws from the set of bear images and the post-change class $\mathbb{P}_{1}$ consists of i.i.d. draws from the set of beaver images (Figure 12). Note that bears and beavers are visually difficult to tell apart, so localization of the changepoint is a difficult task.
@@
We can then use the MCP algorithm and Algorithm 2 to produce a confidence interval. In this case, the left and right p-values are independent, since the left score function $s_{t}^{(0)}$ only depends on data to the left of $t$ and the right score function $s_{t}^{(1)}$ only depends on data to the right of $t$. Therefore, we use the minimum method to combine the p-values, plotting the resulting p-values $(p_{t})_{t=1}^{n-1}$ in Figure 13; the resulting confidence set is $[263,316]$. For a similar experiment regarding a digit change from the MNIST dataset ([deng2012mnist]), see Appendix A.
@@ Acknowledgments
We thank Carlos Padilla for early discussions, Jing Lei for discussions relating to the conformal Neyman-Pearson lemma, Swapnaneel Bhattacharyya for noticing an error in an early preprint, Alessandro Rinaldo for several references to prior work in changepoint localization, and Rohan Hore for providing an implementation of the KCPD algorithm.
Appendix A MNIST digit change
In this section, we consider a simulation of a digit change from the MNIST handwritten digit dataset [deng2012mnist]. Suppose that we observe data of length $n=1000$ and the changepoint is $\xi=400$; the pre-change class $\mathbb{P}_{0}$ consists of i.i.d. draws from the set of handwritten “3” digits and the post-change class $\mathbb{P}_{1}$ consists of i.i.d. draws from the set of handwritten “7” digits (Figure 14).
As described in Section 5, we can use a pretrained multi-class classifier to estimate the likelihood ratio using the data before and after $t$ (for each $t\in[n-1]$). In this case, we use a simple convolutional neural network; see Appendix H for details about how the network was trained.
@@ Appendix B Using certified data to tighten confidence sets
Consider the MNIST digit change experiment of Appendix A but suppose we had additional certified pre-change and post-change samples, which we add to the left and right samples respectively. In this case, we are able to obtain better results.
@@
Figure 16: p-values for an MNIST digit change at $\xi=400$ using 100 certified calibration samples (a) on the left side only and (b) on each side. The dashed red line indicates the true changepoint, and the region on the horizontal axis where the p-values lie above the dotted green line ($\alpha=0.05$) corresponds to our 95% confidence set. The point estimators are (a) $\hat{\xi}=393$ and (b) $\hat{\xi}=405$, both of which are close to the true $\xi=400$.
@@
Appendix C Comparison of combining functions
In this section, we compare empirically between different combining functions. We consider a dataset of size $n=100$ containing a change from $\mathcal{N}(-1,1)$ to $\mathcal{N}(1,1)$ at $\xi=40$. We apply the MCP algorithm to this dataset using the minimum, Fisher, and Bonferroni combining methods outlined in Section 4.2 with the oracle and KDE score functions used in Section 7.1, in order to construct a 95% confidence set. Since the KDE score function is adaptive and thereby induces dependence between the left and right p-values, our theoretical results are technically only valid if one uses the Bonferroni combining method. However, we observe that in practice, the choice of combining method has very little effect on the resulting coverage and width of our confidence sets. We display the results in Table 4.
Method Score function Avg. width Empirical coverage Bias Mean abs. dev.
Minimum Oracle $30.8_{\mathrm{se:}\,10.5}$ $0.95_{\pm 0.03}$ $0.5_{\mathrm{se:}\,6.8}$ $4.9_{\mathrm{se:}\,4.7}$
KDE $33.1_{\mathrm{se:}\,12.3}$ $0.95_{\pm 0.03}$ $1.4_{\mathrm{se:}\,14.8}$ $7.7_{\mathrm{se:}\,12.7}$
Fisher Oracle $31.9_{\mathrm{se:}\,11.3}$ $0.95_{\pm 0.03}$ $0.5_{\mathrm{se:}\,6.1}$ $4.5_{\mathrm{se:}\,4.2}$
KDE $32.9_{\mathrm{se:}\,12.7}$ $0.95_{\pm 0.03}$ $0.1_{\mathrm{se:}\,13.9}$ $7.5_{\mathrm{se:}\,11.7}$
Bonferroni Oracle $31.4_{\mathrm{se:}\,10.7}$ $0.95_{\pm 0.03}$ $2.8_{\mathrm{se:}\,6.8}$ $5.7_{\mathrm{se:}\,4.6}$
KDE $33.6_{\mathrm{se:}\,13.0}$ $0.95_{\pm 0.03}$ $1.0_{\mathrm{se:}\,14.4}$ $7.7_{\mathrm{se:}\,12.2}$
Table 4: Average width and coverage of confidence sets for Gaussian mean change, as well as the bias ($\mathbb{E}[\hat{\xi}-\xi]$) and mean absolute deviation ($\mathbb{E}[\lvert\hat{\xi}-\xi\rvert$]) of the point estimator over 1000 trials. The length of the data is $n=100$ and we observe a change from $\mathcal{N}(-1,1)$ to $\mathcal{N}(1,1)$ at $\xi=40$. We run the MCP algorithm using the oracle and KDE score functions (Section 4.2) to construct a 95% confidence set for the changepoint. Note that the choice of combining function neither significantly affects the width and coverage of the resulting confidence sets, nor the bias and mean absolute deviation of our point estimator. Binomial errors are reported for empirical coverage (95%) and standard errors are reported for all other values.
Appendix D Extensions of the MCP algorithm
@@ D.1 Alternative tests
We can use the asymptotic distribution of $W_{t}$ to construct confidence sets for the changepoint. This method is valid when $t$ and ntn-t are both large. Recall that $\varphi_{t}=B_{t}-tB_{1}$ (where $(B_{t})_{t\geq 0}$ is a standard Brownian motion on $\mathbb{R}$) is the Brownian bridge on $\mathbb{R}$. Define $F_{\varphi}(z)\coloneqq\mathbb{P}\left(\sup_{t\in[0,1]}\;\lvert\varphi_{t}\rvert\leq z\right)$, where $(\varphi_{t})_{t\geq 0}$ is the Brownian bridge. Then, by a central limit theorem due to [donsker1951invariance], we can compute the asymptotically valid p-values
@@
Note that in general, one can use the approximation $p_{t}^{\mathrm{left}}\approx 2\exp(-2\,(W_{t}^{(0)})^{2})$, which is obtained by truncating the series expansion of the limiting distribution ([kolmogorov1933sulla]). In practice, we can use the empirical test to test $\mathcal{H}_{0t}$ when $t$ and ntn-t are small, and use the asymptotic test when $t$ and ntn-t are large. Next, we will discuss how to combine the left and right p-values from Section 4.1 into a single p-value for $\mathcal{H}_{0t}$.
@@
Theorem D.1 (Asymptotic coverage guarantee (asymptotic KS test)).
@@
Input: $(W_{t}^{(0)},W_{t}^{(1)})$ (discrepancy scores), $B\in\mathbb{N}$ (user-chosen sample size)
@@
for $b\in[B]$ do
 $(X_{t}^{(b)})_{t=1}^{n}\sim\mathrm{Unif}([0,1]^{n})$
  Compute $(W_{t}^{(0,b)})_{t=1}^{n-1}$ and $(W_{t}^{(1,b)})_{t=1}^{n-1}$ using Algorithm 1 to Algorithm 1 of Algorithm 1 on permuted data $(X_{t}^{(b)})_{t=1}^{n}$
end for
Draw $\theta_{0},\,\theta_{1}\sim\mathrm{Unif}(0,1)$
$p_{t}^{\mathrm{left}}\leftarrow\frac{1}{B+1}\left(\theta_{0}+\sum_{b=1}^{B}(\mathbf{1}_{W_{t}^{(0,b)}>W_{t}^{(0)}}+\theta_{0}\,\mathbf{1}_{W_{t}^{(0,b)}=W_{t}^{(0)}})\right)$
$p_{t}^{\mathrm{right}}\leftarrow\frac{1}{B+1}\left(\theta_{1}+\sum_{b=1}^{B}(\mathbf{1}_{W_{t}^{(1,b)}>W_{t}^{(1)}}+\theta_{1}\,\mathbf{1}_{W_{t}^{(1,b)}=W_{t}^{(1)}})\right)$
@@
D.2 Alternatives to the Kolmogorov-Smirnov statistic
@@
Appendix E Confidence sets with no changepoint
When there is no changepoint, Algorithm 1 tends to produce very wide confidence sets, but which typically include $n$ (meaning that the practitioner can correctly conclude that there is a possibility that there is no changepoint in the dataset). We summarize these findings in Table 5.
$\alpha$ Avg. width Empirical coverage
0.05 $476.0_{\mathrm{se:}\,53.9}$ $0.95_{\pm 0.03}$
0.5 $250.4_{\mathrm{se:}\,136.5}$ $0.51_{\pm 0.03}$
Table 5: Average width and coverage of confidence sets for Gaussian mean change when there is no changepoint. The length of the data is $n=500$ and all data is from $\mathcal{N}(-1,1)$. We run the MCP algorithm using the oracle score function (Section 4.2) assuming that there is a change from $\mathcal{N}(-1,1)$ to $\mathcal{N}(1,1)$ in order to construct a $1-\alpha$ confidence set for the changepoint. We note that the resulting confidence sets are wide, but typically have the correct coverage. Binomial errors are reported for empirical coverage and standard errors are reported for the width.
Appendix F Proofs of main results
In this section, we present proofs of our main results. For an alternative proof technique, see [vovk2003testing, angelopoulos2024theoretical]. We begin with a useful lemma (for more on this lemma, see [brockwell2007universal]).
Lemma F.1 (Randomized probability integral transform).